# Adversarial Attacks on Data Attribution

### v3

@@ 3.2 The adversary
Assume that the Adversary has access to black-box query to the model $\mathcal{T}(Z_{0})$ to query the model predictions on any input feature $x\in\mathcal{X}$.
@@ 6 Experiments
In this section, we evaluate the effectiveness of the proposed attack methods by the increase of compensation share after manipulating the data with the proposed attack methods.
@@ 6.1 Experimental setup
We consider two machine-learning tasks: image classification and text generation. For image classification, we experiment on MNIST (LeCun, 1998), Digits (Jiang et al., 2023) and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, with different target models including Logistic Regression (LR), Multi-layer Perceptron (MLP), Convolutional Neural Networks (CNN), and ResNet-18 (He et al., 2016). We also employ three popular data attribution methods, including Influence Function (Koh & Liang, 2017), TRAK (Park et al., 2023), and Data Shapley (Ghorbani & Zou, 2019). For text generation, we conduct experiments on NanoGPT (Karpathy, 2022) trained on the Shakespeare dataset (Karpathy, 2015), with TRAK as the data attribution method. Finally, for the image classification settings, we evaluate both Shadow Attack and Outlier Attack, while for the text generation setting, we evaluate Outlier Attack only as 2 usually does not hold for generative AI settings. For the data attribution algorithms, we adopt the implementation from the dattri library (Deng et al., 2024). The experimental settings are summarized in Table 1.
@@ B.3 Attack methods
  • Zeroth Order Optimization (ZOO): ZOO is a black-box attack method that approximates gradient through finite numerical methods. In our experiment, given an input $z=(x,y)$, we perturb it by $x^{\prime}\leftarrow x+\epsilon\cdot\operatorname{sgn}(g(\hat{\theta},z))$, where $g(\hat{\theta},z)$ is an estimation of the loss of the gradient with respect to $x$, given by the symmetric difference quotient
@@
    We modify the method as follows. Firstly, for a character-level GPT model (in our case, NanoGPT), we work on characters rather than words. Secondly, as we work on text generation task, for a sequence of characters that we want to perturb, we take the sum of the negative log-likelihood of predicting its $m$ future characters as a loss. We then sort the characters by their importance, where importance is measured by the loss increase after the character is masked. Finally, the $k$-most important characters are replaced by characters that maximize the increment of the loss. In experiments we set $m=20$ and $k=15$.
@@ C.4 Adversarial attack under data augmentation
In this section, we test the performance of our proposed attack methods when the AI Developer utilize data augmentation techniques when training the model. Specifically, after the AI Developer gathers data, the training dataset is then formed by a combination of the gathered data and its augmented version. For simplicity, we consider settings (a), (c), and (d) in Table 1 and consider standard image augmentation methods such as random cropping and random affine transformation to images. With data augmentation, the compensation share will be attributed back to the original data point if its augmented version is identified as influential. The results are shown in Tables 10 and 11. Compared to the original setting without data augmentation, we cannot draw a definite conclusion on whether data augmentation helps defend the proposed attacks since the trend is unclear. However, overall, we can conclude that the proposed attacks are still highly effective even when data augmentation is utilized by the AI Developer.
@@ C.5 Compensating only correct prediction
In this section we consider the setup when the AI Developer only rewards training data points that are highly influential to validation samples with correct predictions. In detail, after training the model and before calculating of compensation share, the AI Developer will test the model on a private validation set, select those validation points that are correctly predicted, and only attribute compensation to influential training points for these validation points. We test this on settings (a), (c), and (d). Note that the Fraction of Change metric is no longer feasible in this setup since the model trained on the original and manipulated training set may not share the same validation set. The results are shown in Table 12. Overall, our proposed attacks are still highly effective based on the Ratio of Compensation Share.
@@ C.6 Adversarial attack on large-scale setup
In this section, we scale up our experiment and see how the proposed attacks generalize in this large-scale experiment. Specifically, we consider the ResNet-18 model (He et al., 2016) on the Tiny ImageNet dataset (Le & Yang, 2015) with $|Z_{0}|,|Z_{1}|)=(50000,60000)$, and we perturb $|Z_{1}^{a}|=100$ training points. Both the Shadow Attack and Outlier Attack are tested, and the results are shown in Table 13. It is evident that from the Ratio of Compensation Share, two proposed attacks are still highly effective in a large-scale experiment.

### v4

@@ 3.2 The adversary
Assume that the Adversary has access to a black-box query to the model $\mathcal{T}(Z_{0})$ to query the model predictions on any input feature $x\in\mathcal{X}$.
@@ 6 Experiments
In this section, we evaluate the effectiveness of the proposed attack methods by the increase in compensation share after manipulating the data with the proposed attack methods.
@@ 6.1 Experimental setup
We consider two machine-learning tasks: image classification and text generation. For image classification, we experiment on MNIST (LeCun, 1998), Digits (Jiang et al., 2023), and CIFAR-10 (Krizhevsky & Hinton, 2009) datasets, with different target models including Logistic Regression (LR), Multi-layer Perceptron (MLP), Convolutional Neural Networks (CNN), and ResNet-18 (He et al., 2016). We also employ three popular data attribution methods, including Influence Function (Koh & Liang, 2017), TRAK (Park et al., 2023), and Data Shapley (Ghorbani & Zou, 2019). For text generation, we conduct experiments on NanoGPT (Karpathy, 2022) trained on the Shakespeare dataset (Karpathy, 2015), with TRAK as the data attribution method. Finally, for the image classification settings, we evaluate both Shadow Attack and Outlier Attack, while for the text generation setting, we evaluate Outlier Attack only, as 2 usually does not hold for generative AI settings. For the data attribution algorithms, we adopt the implementation from the dattri library (Deng et al., 2024). The experimental settings are summarized in Table 1.
@@ B.3 Attack methods
  • Zeroth Order Optimization (ZOO): ZOO is a black-box attack method that approximates the gradient through finite numerical methods. In our experiment, given an input $z=(x,y)$, we perturb it by $x^{\prime}\leftarrow x+\epsilon\cdot\operatorname{sgn}(g(\hat{\theta},z))$, where $g(\hat{\theta},z)$ is an estimation of the loss of the gradient with respect to $x$, given by the symmetric difference quotient
@@
    We modify the method as follows. Firstly, for a character-level GPT model (in our case, NanoGPT), we work on characters rather than words. Secondly, as we work on the text generation task, for a sequence of characters that we want to perturb, we take the sum of the negative log-likelihood of predicting its $m$ future characters as a loss. We then sort the characters by their importance, where importance is measured by the loss increase after the character is masked. Finally, the $k$-most important characters are replaced by characters that maximize the increment of the loss. In experiments we set $m=20$ and $k=15$.
@@ C.4 Adversarial attack under data augmentation
In this section, we test the performance of our proposed attack methods when the AI Developer utilizes data augmentation techniques when training the model. Specifically, after the AI Developer gathers data, the training dataset is then formed by a combination of the gathered data and its augmented version. For simplicity, we consider settings (a), (c), and (d) in Table 1 and consider standard image augmentation methods such as random cropping and random affine transformation of images. With data augmentation, the compensation share will be attributed back to the original data point if its augmented version is identified as influential. The results are shown in Tables 10 and 11. Compared to the original setting without data augmentation, we cannot draw a definite conclusion on whether data augmentation helps defend against the proposed attacks since the trend is unclear. However, overall, we can conclude that the proposed attacks are still highly effective even when data augmentation is utilized by the AI Developer.
@@ C.5 Compensating only correct prediction
In this section, we consider the setup when the AI Developer only rewards training data points that are highly influential to validation samples with correct predictions. In detail, after training the model and before calculating of compensation share, the AI Developer will test the model on a private validation set, select those validation points that are correctly predicted, and only attribute compensation to influential training points for these validation points. We test this on settings (a), (c), and (d). Note that the Fraction of Change metric is no longer feasible in this setup since the model trained on the original and manipulated training set may not share the same validation set. The results are shown in Table 12. Overall, our proposed attacks are still highly effective based on the Ratio of Compensation Share.
@@ C.6 Adversarial attack on large-scale setup
In this section, we scale up our experiment and see how the proposed attacks generalize in this large-scale experiment. Specifically, we consider the ResNet-18 model (He et al., 2016) on the Tiny ImageNet dataset (Le & Yang, 2015) with $|Z_{0}|,|Z_{1}|)=(50000,60000)$, and we perturb $|Z_{1}^{a}|=100$ training points. Both the Shadow Attack and Outlier Attack are tested, and the results are shown in Table 13. It is evident that from the Ratio of Compensation Share, the two proposed attacks are still highly effective in a large-scale experiment.